OpenStudy (anonymous):

A zucchini plant in Darnell’s garden was 9 centimeters tall when it was first planted. Since then, it has grown approximately 0.5 centimeters per day. a. Write a rule to describe the function. b. After how many days will the zucchini plant be 0.185 meters tall? (Hint: 1 meter = 100 centimeters)

OpenStudy (gschibby):

f(x) = 0.5x + 9

OpenStudy (gschibby):

Then you set f(x) = 18.5 and solve for x ;)

OpenStudy (gschibby):

\[0.5x+9 = 18.5\] \[x = \frac{ 18.5 - 9 }{ 0.5 } = 19\] So 19 days ;)